OpenStudy (anonymous):

The probability that a dessert sold at a certain café contains chocolate is 73%. The probability that a dessert contains both chocolate and nuts is 25%. Find the probability that a randomly chosen dessert contains nuts. Use the formula P(B|A) = P(A and B). (1 point) P(A) a. 29.2% b. 48.0% c. 34.2% d. 36.1%

ganeshie8 (ganeshie8):

I think they want you find below : Find the probability that a randomly chosen dessert contains nuts given that it contains chocolate.

ganeshie8 (ganeshie8):

then u can do : P(nuts | chocolate) = P(nuts AND chocolate)/P(chocolate) = 25/73 = 34.2%